OpenStudy (anonymous):

compute the integrals ∫ x/√1-2x dx

OpenStudy (zarkon):

\[\int\frac{x}{\sqrt{1-2x}}dx\]?

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

let \(u=1-2x\)

OpenStudy (anonymous):

is du= -2 dx ??

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

so is it rght -1/2 ∫ du/ u ?

OpenStudy (zarkon):

no

OpenStudy (zarkon):

solve \[u=1-2x\] for x and use that to replace the x in the numerator